OpenStudy (anonymous):

Perimeter of a Square Perimeter of a Rectangle Area of a Square Area of a Rectangle Area of a Triangle No Match A.4S B.s^2 C.L+W D.L*W E. b*h/2 F.2L+2W

OpenStudy (anonymous):

@suyogsomavanshi

OpenStudy (anonymous):

Perimeter of a Square = 4s Perimeter of a Rectangle = 2L+2W Area of a Square = s^2 Area of a Rectangle =L*W Area of a Triangle= b*h/2 Is it fine ???

OpenStudy (anonymous):

thnkz I was only confused with 2.

OpenStudy (anonymous):

got it now.

OpenStudy (anonymous):

You are welcome :) bye :)